TheKayla:

What is the answer to the following: 1/2 + x =3

UkuleleGirl:

Subtract \[\frac{ 1 }{ 2 }\] from both sides first, then simplify

UkuleleGirl:

you would get \[\frac{ 5 }{ 2 }\]

UkuleleGirl:

Can you tell me what you simplify that to?

mdobbs6856:

1/2+x=3 subtract 1/2 from both sides x=3-1/2 then you'll get 5/2 or 2 1/2 or 2.5

TheKayla:

@mdobbs6856 wrote:
1/2+x=3 subtract 1/2 from both sides x=3-1/2 then you'll get 5/2 or 2 1/2 or 2.5
Thank you!

TheKayla:

@ukulelegirl wrote:
Can you tell me what you simplify that to?
I just needed the fractional answer. Not simplification.

mdobbs6856:

@thekayla wrote:
@mdobbs6856 wrote:
1/2+x=3 subtract 1/2 from both sides x=3-1/2 then you'll get 5/2 or 2 1/2 or 2.5
Thank you!
Np!!!!

UkuleleGirl:

OKay so you simplify it to a decimal.